# Off-Road Navigation via Implicit Neural Representation of Terrain Traversability

### v1

@@ III Simulation Experiments
We evaluate the proposed approach on a series of representative off-road navigation scenarios in a Unity simulation environment.
@@ III-E Navigation Performance
Path 1 Geometry 100.0 $\textbf{100.0}\pm\textbf{0.0}$ $\underline{44.0\pm 1.6}$ $\underline{24.0\pm 1.5}$ $\underline{1.8\pm 0.1}$ $\underline{6.1\pm 0.3}$
@@
TRAIL 100.0 $\textbf{100.0}\pm\textbf{0.0}$ $\textbf{11.5}\pm\textbf{2.2}$ $\textbf{11.4}\pm\textbf{2.2}$ $\textbf{1.7}\pm\textbf{0.1}$ $\textbf{6.0}\pm\textbf{1.7}$
Path 2 Geometry 100.0 $\textbf{100.0}\pm\textbf{0.0}$ $\underline{38.3\pm 14.2}$ $\underline{20.9\pm 9.0}$ $\underline{1.8\pm 0.1}$ $\underline{6.2\pm 0.1}$
@@
TRAIL 100.0 $\textbf{100.0}\pm\textbf{0.0}$ $\textbf{10.9}\pm\textbf{1.1}$ $\textbf{10.8}\pm\textbf{1.0}$ $\textbf{1.7}\pm\textbf{0.1}$ $\textbf{5.1}\pm\textbf{0.7}$
@@
Path 1 Geometry 100.0 $\textbf{100.0}\pm\textbf{0.0}$ $26.8\pm 6.7$ $14.2\pm 1.3$ $1.9\pm 0.2$ $\textbf{5.5}\pm\textbf{0.5}$
WayFASTER 100.0 $\textbf{100.0}\pm\textbf{0.0}$ $\underline{11.5\pm 1.8}$ $\underline{10.9\pm 0.9}$ $\underline{1.8\pm 0.2}$ $\underline{5.5\pm 0.8}$
TRAIL 100.0 $\textbf{100.0}\pm\textbf{0.0}$ $\textbf{10.9}\pm\textbf{0.4}$ $\textbf{10.9}\pm\textbf{0.4}$ $\textbf{1.8}\pm\textbf{0.1}$ $6.0\pm 0.3$
@@
WayFASTER 100.0 $\textbf{100.0}\pm\textbf{0.0}$ $\underline{13.3\pm 1.7}$ $\textbf{12.1}\pm\textbf{0.8}$ $\underline{1.9\pm 0.1}$ $\underline{5.8\pm 0.2}$
TRAIL 100.0 $\textbf{100.0}\pm\textbf{0.0}$ $\textbf{12.3}\pm\textbf{0.7}$ $\underline{12.5\pm 1.0}$ $\textbf{1.7}\pm\textbf{0.1}$ $\textbf{5.5}\pm\textbf{0.5}$
@@
Path 1 Geometry 100.0 $\textbf{100.0}\pm\textbf{0.0}$ $52.9\pm 10.4$ $19.1\pm 3.0$ $\textbf{1.7}\pm\textbf{0.1}$ $14.3\pm 12.7$
WayFASTER 100.0 $\textbf{100.0}\pm\textbf{0.0}$ $\textbf{10.8}\pm\textbf{0.1}$ $\textbf{10.4}\pm\textbf{0.1}$ $1.8\pm 0.0$ $\underline{5.7\pm 0.4}$
TRAIL 100.0 $\textbf{100.0}\pm\textbf{0.0}$ $\underline{10.9\pm 0.3}$ $\underline{11.0\pm 0.2}$ $\textbf{1.7}\pm\textbf{0.1}$ $\textbf{4.5}\pm\textbf{0.7}$
Path 2 Geometry $0.0$ $\underline{67.0\pm 1.2}$ N/A N/A N/A N/A
@@
TRAIL 66.7 $\textbf{91.5}\pm\textbf{12.1}$ $\textbf{12.9}\pm\textbf{0.5}$ $\textbf{11.6}\pm\textbf{0.1}$ $\textbf{2.0}\pm\textbf{0.4}$ $\textbf{31.6}\pm\textbf{37.0}$
@@
Path 1 Geometry $\underline{66.7}$ $\underline{79.1\pm 29.5}$ $\underline{33.5\pm 11.8}$ $\underline{13.7\pm 1.3}$ $\underline{4.0\pm 1.7}$ $\underline{79.3\pm 54.0}$
@@
TRAIL 100.0 $\textbf{100.0}\pm\textbf{0.0}$ $\textbf{12.7}\pm\textbf{0.2}$ $\textbf{12.5}\pm\textbf{0.2}$ $\textbf{1.9}\pm\textbf{0.2}$ $\textbf{5.1}\pm\textbf{0.8}$
Path 2 Geometry $0.0$ $\underline{60.4\pm 4.2}$ N/A N/A N/A N/A
@@
TRAIL 100.0 $\textbf{100.0}\pm\textbf{0.0}$ $\textbf{15.3}\pm\textbf{1.9}$ $\textbf{14.7}\pm\textbf{1.3}$ $\textbf{3.6}\pm\textbf{0.3}$ $\textbf{52.2}\pm\textbf{17.8}$
@@ III-F Ablation Study
Time $\mathrm{s}$] $\mathbf{12.0\pm 0.4}$ N/A $17.3\pm 2.2$ $17.2\pm 0.4$} $54.8\pm 4.6$ $35.6\pm 18.4$
Length $\mathrm{m}$] $\mathbf{10.3\pm 0.4}$ N/A $11.1\pm 0.7$ $10.6\pm 0.2$ $15.5\pm 0.6$ $14.2\pm 3.5$
Vert. Acc. RMS $\mathrm{m}\text{\,}{\mathrm{s}}^{-2}$] $\mathbf{1.8\pm 0.3}$ N/A $3.0\pm 0.3$ $2.0\pm 0.3$ $1.9\pm 0.2$ $2.8\pm 0.6$
Vert. Acc. Max $\mathrm{m}\text{\,}{\mathrm{s}}^{-2}$] $\mathbf{4.4\pm 0.5}$ N/A $61.4\pm 26.6$ $21.5\pm 23.7$ $41.6\pm 26.1$ $87.3\pm 9.2$
@@ IV Field Experiments
Figure 7: GPS locations of the robot during the long path navigation experiment. The total length is about $370$\mathrm{m}$\$.
@@
Figure 8: Short Path Navigation Results. Each column shows one set of experiments with separate trials colored differently. The first row shows the paths followed by Geometry, the second row shows the paths followed by WayFaster, and the third row shows the paths followed by TRAIL. The visualizations are obtained by projecting the odometry to the onboard camera image captured at the starting position.

### v2

@@ III Simulation Experiments
We first evaluate the proposed approach on a series of representative off-road navigation scenarios in a Unity simulation environment.
@@ III-E Navigation Performance
Path 1 Geometry 100.0 100.0 $\pm$ 0.0 44.0 $\pm$ 1.6 24.0 $\pm$ 1.5 1.8 $\pm$ 0.1 6.1 $\pm$ 0.3
@@
TRAIL 100.0 100.0 $\pm$ 0.0 11.5 $\pm$ 2.2 11.4 $\pm$ 2.2 1.7 $\pm$ 0.1 6.0 $\pm$ 1.7
Path 2 Geometry 100.0 100.0 $\pm$ 0.0 38.3 $\pm$ 14.2 20.9 $\pm$ 9.0 1.8 $\pm$ 0.1 6.2 $\pm$ 0.1
@@
TRAIL 100.0 100.0 $\pm$ 0.0 10.9 $\pm$ 1.1 10.8 $\pm$ 1.0 1.7 $\pm$ 0.1 5.1 $\pm$ 0.7
@@
Path 1 Geometry 100.0 100.0 $\pm$ 0.0 26.8 $\pm$ 6.7 14.2 $\pm$ 1.3 1.9 $\pm$ 0.2 5.5 $\pm$ 0.5
WayFASTER 100.0 100.0 $\pm$ 0.0 11.5 $\pm$ 1.8 10.9 $\pm$ 0.9 1.8 $\pm$ 0.2 5.5 $\pm$ 0.8
TRAIL 100.0 100.0 $\pm$ 0.0 10.9 $\pm$ 0.4 10.9 $\pm$ 0.4 1.8 $\pm$ 0.1 6.0 $\pm$ 0.3
@@
WayFASTER 100.0 100.0 $\pm$ 0.0 13.3 $\pm$ 1.7 12.1 $\pm$ 0.8 1.9 $\pm$ 0.1 5.8 $\pm$ 0.2
TRAIL 100.0 100.0 $\pm$ 0.0 12.3 $\pm$ 0.7 12.5 $\pm$ 1.0 1.7 $\pm$ 0.1 5.5 $\pm$ 0.5
@@
Path 1 Geometry 100.0 100.0 $\pm$ 0.0 52.9 $\pm$ 10.4 19.1 $\pm$ 3.0 1.7 $\pm$ 0.1 14.3 $\pm$ 12.7
WayFASTER 100.0 100.0 $\pm$ 0.0 10.8 $\pm$ 0.1 10.4 $\pm$ 0.1 1.8 $\pm$ 0.0 5.7 $\pm$ 0.4
TRAIL 100.0 100.0 $\pm$ 0.0 10.9 $\pm$ 0.3 11.0 $\pm$ 0.2 1.7 $\pm$ 0.1 4.5 $\pm$ 0.7
Path 2 Geometry 0.0 67.0 $\pm$ 1.2 N/A N/A N/A N/A
@@
TRAIL 66.7 91.5 $\pm$ 12.1 12.9 $\pm$ 0.5 11.6 $\pm$ 0.1 2.0 $\pm$ 0.4 31.6 $\pm$ 37.0
@@
Path 1 Geometry 66.7 79.1 $\pm$ 29.5 33.5 $\pm$ 11.8 13.7 $\pm$ 1.3 4.0 $\pm$ 1.7 79.3 $\pm$ 54.0
@@
TRAIL 100.0 100.0 $\pm$ 0.0 12.7 $\pm$ 0.2 12.5 $\pm$ 0.2 1.9 $\pm$ 0.2 5.1 $\pm$ 0.8
Path 2 Geometry 0.0 60.4 $\pm$ 4.2 N/A N/A N/A N/A
@@
TRAIL 100.0 100.0 $\pm$ 0.0 15.3 $\pm$ 1.9 14.7 $\pm$ 1.3 3.6 $\pm$ 0.3 52.2 $\pm$ 17.8
@@ III-F Ablation Study
Time $\mathrm{s}$] 12.0 $\pm$ 0.4 N/A 17.3 $\pm$ 2.2 17.2 $\pm$ 0.4 54.8 $\pm$ 4.6 35.6 $\pm$ 18.4
Length $\mathrm{m}$] 10.3 $\pm$ 0.4 N/A 11.1 $\pm$ 0.7 10.6 $\pm$ 0.2 15.5 $\pm$ 0.6 14.2 $\pm$ 3.5
Vert. Acc. RMS $\mathrm{m}\text{\,}{\mathrm{s}}^{-2}$] 1.8 $\pm$ 0.3 N/A 3.0 $\pm$ 0.3 2.0 $\pm$ 0.3 1.9 $\pm$ 0.2 2.8 $\pm$ 0.6
Vert. Acc. Max $\mathrm{m}\text{\,}{\mathrm{s}}^{-2}$] 4.4 $\pm$ 0.5 N/A 61.4 $\pm$ 26.6 21.5 $\pm$ 23.7 41.6 $\pm$ 26.1 87.3 $\pm$ 9.2
@@ IV Field Experiments
Figure 7: GPS locations of the robot during the long path navigation hardware experiment. The total length is about $370$\mathrm{m}$\$.
@@
Figure 8: Short Path Navigation Results. Each column shows one set of experiments with separate trials colored differently. The first row shows the paths followed by Geometry, the second row shows the paths followed by WayFaster, and the third row shows the paths followed by TRAIL. The visualizations are obtained by projecting the odometry to the onboard camera image captured at the starting position. In the fifth experiment (last column), TRAIL was able to find a clearing on the right (paved road) and successfully navigate to the goal, while the two baseline methods were trapped by twigs (which are hard to distinguish from the dry grass).